All Systems Operational

About This Site

This website shows live status & latency information about the DataSift platform.

Frontend ? Operational
PYLON for Facebook Topic Data Operational
Public Sources ? Operational
Managed Sources ? Operational
Augmentations ? Operational
HTTP(s) Streaming ? Operational
Push System ? Operational
Historics ? Operational
Filtering ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Past Incidents
Aug 25, 2016

No incidents reported today.

Aug 24, 2016

No incidents reported.

Aug 23, 2016

No incidents reported.

Aug 22, 2016

No incidents reported.

Aug 21, 2016

No incidents reported.

Aug 20, 2016

No incidents reported.

Aug 19, 2016
Completed - The scheduled maintenance has been completed.
Aug 19, 09:39 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 18, 15:01 UTC
Scheduled - Due to unexpected issues, this maintenance window is to be repeated 2016-08-18 at 15:00 UTC. Our apologies for the inconvenience.
Aug 18, 13:17 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 17, 15:00 UTC
Update - Tumblr data to be unavailable for approximately 1 hour from the start of this maintenance window.
Aug 16, 21:27 UTC
Scheduled - Due to upstream maintenance, Tumblr data will be unavailable for a 1 hour window during this time.

This maintenance notification will be updated once an exact date and time is known.
Aug 5, 15:04 UTC
Aug 17, 2016
Completed - The scheduled maintenance has been completed.
Aug 17, 16:01 UTC
Verifying - DNS maintenance actions are now complete.

We will be performing rolling restarts of the old cluster every 8 hours over the next few days to prompt re-connections to the new load balanced configuration.
Aug 17, 15:49 UTC
Update - The legacy DNS records (46.236.5.226) have now been removed. As TTLs and caches expire all stream.datasift.com / websocket.datasift.com A records will resolve to 185.20.5.45.

We anticipate clients will gradually drain from the old cluster accordingly.
Aug 17, 15:11 UTC
Update - We are currently publishing both DNS A records in a round robin fashion and seeing clients connecting to both clusters.

In the next ~60minutes we will remove the old A record.
Aug 17, 14:26 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 17, 12:00 UTC
Scheduled - At 12:00 UTC we will be migrating the DNS records for stream.datasift.com from IPv4 address 46.236.5.226 to IPv4 address 185.20.5.45

This DNS change will result in customers accessing the streaming service via a new loadbalancer setup.
Aug 10, 10:22 UTC
Completed - The scheduled maintenance has been completed.
Aug 17, 11:04 UTC
Scheduled - We will be performing routing maintenance in one of our POPs to improve redundancy.

No outage is expected, however there will be an at risk period during this maintenance.
Aug 9, 11:26 UTC
Aug 16, 2016

No incidents reported.

Aug 15, 2016

No incidents reported.

Aug 14, 2016

No incidents reported.

Aug 13, 2016

No incidents reported.

Aug 12, 2016

No incidents reported.

Aug 11, 2016

No incidents reported.